Solve the y+ 4y = initial value problem s 1 if 0<xsa To if x>,T ylo)= 1, g(0)=0
Answer #1

Y + 4y = if OS* 911 Y(0) El; y col = 0 0 7 X>7 +(+) - I(u() - u (t-Tl) = u(t) -u (7-77) f(x)} = { -TTS e s Apply Laplace tra